OpenStudy (anonymous):

Evaluate. 6[5(3-9)-1]+2 --------------- 7(8-2)+4

TheSmartOne (thesmartone):

This will help you. http://openstudy.com/study#/updates/545826d0e4b04ccc259160ab

TheSmartOne (thesmartone):

\(\Huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

@dudezack @~

OpenStudy (anonymous):

Simplify. 6[5(3-9)-1]+2 --------------- 7(8-2)+4

OpenStudy (michele_laino):

@dudezack please solve th parenthesis (3-9) first

OpenStudy (anonymous):

-6

OpenStudy (michele_laino):

ok, now multiply -6 by 5

OpenStudy (anonymous):

-30

OpenStudy (michele_laino):

ok, now add -1 to -30

OpenStudy (anonymous):

-62

OpenStudy (michele_laino):

it is -31

OpenStudy (anonymous):

-31

OpenStudy (michele_laino):

ok, now multiply -31 by 5

OpenStudy (anonymous):

-155

OpenStudy (michele_laino):

sorry mltipluy -31 by 6

OpenStudy (anonymous):

-186

OpenStudy (michele_laino):

perfect, now add +2 to -186

OpenStudy (anonymous):

is + a positive

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

please you have to calculate -186+2=...

OpenStudy (anonymous):

-184

OpenStudy (michele_laino):

well done, that is your numerator. Now let us think to denominator

OpenStudy (michele_laino):

please solve the parenthesis (8-2)

OpenStudy (anonymous):

6

OpenStudy (michele_laino):

that's right, now please multiply 6 by 7

OpenStudy (anonymous):

42

OpenStudy (michele_laino):

ok! Now add 4 to 42

OpenStudy (anonymous):

46

OpenStudy (michele_laino):

perfect! that is your denominator, so your fraction is: \[-\frac{ 186 }{ 46 }\]

OpenStudy (anonymous):

how do we didvide it

OpenStudy (michele_laino):

oops... sorry: \[-\frac{ 184 }{ 46 }\]

OpenStudy (michele_laino):

please you have to find the maximum common divisor. in order to do that, note that: 46=2*23 23 is a prime number, and 184=2^3*23 so by definition what is the maximum common divisor?

OpenStudy (anonymous):

idk

OpenStudy (michele_laino):

@dudezack do you know how to calculate the maximum common divisor?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

the rule to find it is: you have to pick only common factors whose exponent is the littlest

OpenStudy (michele_laino):

nemely common factors between 184 and 46 are?

OpenStudy (michele_laino):

@dudezack please compare the factorization above

OpenStudy (anonymous):

ok thanks

OpenStudy (michele_laino):

write your answer, remember you have to pick only common factors

OpenStudy (michele_laino):

for example 2 is a common factor? @dudezack

OpenStudy (michele_laino):

yes it is, because it is a factor of 46, and it is also a factor of 184

OpenStudy (michele_laino):

do you agree @dudezack

OpenStudy (michele_laino):

@dudezack another common factor is 23, do you agree?